OpenStudy (anonymous):

When you calculate a mean, the sign of the quotient always depend on the sign of the dividend. Why?

OpenStudy (anonymous):

This is because the mean is computed by adding two numbers and dividing by 2. You are always dividing by a positive number, so it will only depend on the sign of the result in the dividend. Examples: (-2 + -4) / 2 = -6 / 2 = -6 (-6 means negative result) (2 + 6) / 2 = 8 / 2 = 4 (+8 means positive result) (-7 + 9) / 2 = 2 / 2 = 1 (+2 means positive result)